A ball is released from a window that is 10 meters above the ground. When the ball leaves your hand, it is moving 8m/s at an angle 300 below the horizontal. Ignoring the effect of air resistance, how far horizontally from the window will the ball hit the ground?
1
Expert's answer
2021-08-23T04:49:45-0400

"U_x=ucosα=8*cos30=6.92820m/s\\

U_y=usinα=8*sin30 =4m/s\\

Y=usinα*t+\frac{1}{2}gt^2\\

10=4t+6.92820t^2\\

Solving \space for \space t=0.95sec\\

Horizontal \space distance=U_x*t=6.92820*0.95=6.58179m"